Given that N = {1, 2, 3, …, 100}, then

(i) Write the subset A of N, whose elements are odd numbers.

(ii) Write the subset B of N, whose elements are represented by x + 2, where x ∈ N

Grade:12

1 Answers

Harshit Singh
3 years ago
Welcome to AskIITians

(i) A = {x | x ∈ N and x is odd}

A = {1, 3, 5, 7, …, 99}

(ii) B = {y | y = x + 2, x ∈ N}

1 ∈ N, y = 1 + 2 = 3

2 ∈ N, y = 2 + 2 = 4, and so on.

Therefore, B = {3, 4, 5, 6, … , 100}

Thanks